# Adversary-Aware Private Inference over Wireless Channels

### v1

@@ IV Reconstruction Attack Analysis
Remark 1.
We adopt Laplacian-distributed encoder weights instead of the more conventional Gaussian assumption for two main reasons. First, empirical studies have shown that weight distributions in trained neural networks often exhibit heavy-tailed, Laplacian-like behavior rather than Gaussian, especially in early or sparsely regularized layers [22, 23]. Second, Laplacian distributions yield sharper concentration bounds on the spectral norm [24], which is crucial in our theoretical analysis for bounding privacy leakage and adversarial reconstruction error.
@@
Theorem 2 (Lower Bound on Adversarial MSE for the raw data $\mathbf{x}$).
Let $\hat{\mathbf{z}}$ be the adversary’s MMSE estimator of $\mathbf{z}$. Assume the adversary attempts to recover the feature representation $\mathbf{f}(\mathbf{x})\in\mathbb{R}^{d}$ using the pseudoinverse; i.e., the best linear unbiased estimator:
$\hat{\mathbf{f}}(\mathbf{x})=\mathbf{W}^{\dagger}\hat{\mathbf{z}},\quad\text{where }\mathbf{W}^{\dagger}=(\mathbf{W}^{\top}\mathbf{W})^{-1}\mathbf{W}^{\top}.$
Then the error in reconstructing $\mathbf{f}(\mathbf{x})$ satisfies
$\hat{\mathbf{f}}(\mathbf{x})-\mathbf{f}(\mathbf{x})=\mathbf{W}^{\dagger}(\hat{\mathbf{z}}-\mathbf{z}),$
and the MSE satisfies
$\mathbb{E}\left[\|\hat{\mathbf{f}}(\mathbf{x})-\mathbf{f}(\mathbf{x})\|_{2}^{2}\right]$ $\leq\|\mathbf{W}^{\dagger}\|_{\mathrm{op}}^{2}\cdot\mathbb{E}\left[\|\hat{\mathbf{z}}-\mathbf{z}\|_{2}^{2}\right]$
$\leq\|\mathbf{W}^{\dagger}\|_{\mathrm{op}}^{2}\cdot\mathrm{MSE}_{\mathrm{adv}}.$ (15)
Now suppose $\mathbf{W}$ is a random matrix with i.i.d. entries drawn from Laplacian distribution. Then with high probability over the draw of $\mathbf{W}$,
$\|\mathbf{W}^{\dagger}\|_{\mathrm{op}}\leq\frac{1}{c(\sqrt{d}-\sqrt{r}-t)},\quad\text{for any }t>0,$
with probability at least $1-2e^{-ct^{2}}$. Thus
$\mathbb{E}\left[\|\hat{\mathbf{f}}(\mathbf{x})-\mathbf{f}(\mathbf{x})\|_{2}^{2}\right]\geq\frac{1}{c^{2}(\sqrt{d}-\sqrt{r}-t)^{2}}\cdot\mathrm{MSE}_{\mathrm{adv}}.$
@@ V A Wireless Data Acquisition Model
$\mathbb{E}\left[\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}^{2}\right]\geq\frac{1}{c^{2}(\sqrt{d}-\sqrt{r}-t)^{2}}\cdot\mathrm{MSE}_{\mathrm{adv}},$

### v2

@@ IV Reconstruction Attack Analysis
Theorem 2 (Lower bound on adversarial MSE for recovering $\mathbf{f}(\mathbf{x})$).
Let $\hat{\mathbf{z}}$ be the adversary’s MMSE estimator of $\mathbf{z}$, and suppose the adversary forms
$\hat{\mathbf{f}}(\mathbf{x})=\mathbf{W}^{\dagger}\hat{\mathbf{z}},\qquad\mathbf{W}^{\dagger}:=(\mathbf{W}^{\top}\mathbf{W})^{-1}\mathbf{W}^{\top},$
where $\mathbf{W}\in\mathbb{R}^{r\times d}$ has full row rank (so $d\geq r$). Let $\mathbf{e}:=\hat{\mathbf{z}}-\mathbf{z}$. Then
$\hat{\mathbf{f}}(\mathbf{x})-\mathbf{f}(\mathbf{x})=\mathbf{W}^{\dagger}\mathbf{e}.$
Conditioning on $\mathbf{W}$, we have
$\mathbb{E}\!\left[\|\hat{\mathbf{f}}(\mathbf{x})-\mathbf{f}(\mathbf{x})\|_{2}^{2}\mid\mathbf{W}\right]$ $=\mathbb{E}\!\left[\|\mathbf{W}^{\dagger}\mathbf{e}\|_{2}^{2}\right]$
$\geq\sigma_{\min}^{2}(\mathbf{W}^{\dagger})\,\cdot\mathbb{E}\!\left[\|\mathbf{e}\|_{2}^{2}\right]$
$=\frac{1}{\sigma_{\max}^{2}(\mathbf{W})}\,\cdot\mathrm{MSE}_{\mathrm{adv}}.$ (15)
Now assume $\mathbf{W}$ has i.i.d. zero-mean, unit-variance Laplace entries (more generally, i.i.d. isotropic sub-exponential entries). Then there exist absolute constants $c_{0},c>0$ such that for any $t>0$, with probability at least $1-2e^{-ct^{2}}$,
$\sigma_{\max}(\mathbf{W})\leq c_{0}\big(\sqrt{d}+\sqrt{r}+t\big).$ (16)
Combining (15) and (16) yields: with probability at least $1-2e^{-ct^{2}}$,
$\mathbb{E}\!\left[\|\hat{\mathbf{f}}(\mathbf{x})-\mathbf{f}(\mathbf{x})\|_{2}^{2}\right]\geq\frac{1}{c_{0}^{2}\big(\sqrt{d}+\sqrt{r}+t\big)^{2}}\,\cdot\mathrm{MSE}_{\mathrm{adv}}.$
@@ V A Wireless Data Acquisition Model
$\mathbb{E}\left[\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}^{2}\right]\geq\frac{1}{c_{0}^{2}(\sqrt{d}+\sqrt{r}+t)^{2}}\cdot\mathrm{MSE}_{\mathrm{adv}},$